# Unveiling the jet angular broadening with photon-tagged jets in high-energy nuclear collisions

### v2

@@ Abstract
The medium modification of jet substructure in hot and dense nuclear matter has garnered significant interest from the heavy-ion physics community in recent years. Measurements of inclusive jets show an angular narrowing in nucleus-nucleus collisions, while recent CMS results for photon-tagged jets ($\gamma$+jets) suggest evidence of broadening. In this study, we conduct a theoretical analysis of the angular structure of inclusive jets and $\gamma$+jets using a transport approach that accounts for jet energy loss and the medium response in the quark-gluon plasma. We examine the girth modification of $\gamma$+jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, achieving satisfactory agreement with recent CMS measurements. We explore the relationship between selection bias and jet kinematics by varying the threshold for $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. Notably, we quantitatively demonstrate that $\gamma$+jets significantly reduce selection bias and can effectively select jets that have been sufficiently quenched in PbPb collisions, which is crucial for capture the jet angular broadening. Additionally, we estimate the contributions of medium-induced gluon radiation and the medium response to the broadening of the jet angular substructure. Lastly, we analyze the modification patterns of jet $R_{g}$ and $\Delta R_{\rm axis}$ in PbPb collisions, which indicate slight broadening for $\gamma$+jets and noticeable narrowing for inclusive jets compared to pp collisions.
@@ I Introduction
High-energy collisions of heavy nuclei at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) provide an experimental avenue to unravel the mysteries of the quark-gluon plasma (QGP), a short-lived state of de-confined nuclear matter created at extremely high temperature and density. The jet quenching phenomenon, energy dissipation of an energetic parton when passing through the hot and dense nuclear matter, is one of the most important signatures of the QGP formation Gyulassy:2003mc ; Gyulassy:1990ye ; Qin:2015srf ; Vitev:2008rz ; Proceedings:2007ctk ; Vitev:2009rd ; Casalderrey-Solana:2014bpa ; Gyulassy:1993hr ; Wang:2001ifa ; Vitev:2008vk . Investigations on jet quenching reveal the phase structure of the strongly-coupled nuclear matter and push our knowledge of the quantum chromodynamics (QCD) under extreme conditions Connors:2017ptx ; Andrews:2018jcm ; Cao:2020wlm ; Cunqueiro:2021wls ; Apolinario:2022vzg ; Sorensen:2023zkk ; Zhao:2020wcd ; Xie:2024xbn ; Zhang:2021xib ; Tang:2020ame ; Shen:2020mgh ; Chen:2024eaq ; Chen:2024aom ; Zhao:2022dac ; Shou:2024uga .
@@
Jet substructures are valuable tools to gain insight into the details of the jet-medium interaction in the QGP, such as the medium-induced gluon radiation Baier:1996kr ; Baier:1996sk , medium response  KunnawalkamElayavalli:2017hxo ; Pablos:2019ngg ; Chen:2020tbl ; Casalderrey-Solana:2020rsj ; He:2018xjv ; Ke:2020clc , medium resolution length Hulcher:2017cpt ; Mehtar-Tani:2016aco ; Caucal:2018dla and the “Molière elastic scattering” DEramo:2012uzl ; DEramo:2018eoy . Recent reviews on this topic can be found in references Cunqueiro:2021wls ; CMS:2024krd ; Arslandok:2023utm ; Apolinario:2022vzg ; Marzani:2019hun . A significant focus of recent investigations has been on how the angular structure of jets is modified in nucleus-nucleus collisions—whether they narrow or broaden—an issue that has garnered much attention Ringer:2019rfk ; Rajagopal:2016uip ; Chien:2016led ; Mehtar-Tani:2016aco ; Larkoski:2017bvj ; Chang:2017gkt ; Casalderrey-Solana:2016jvj ; Kang:2023ycg ; Kang:2023qxb ; JETSCAPE:2023hqn ; Milhano:2017nzm ; Caucal:2019uvr ; Casalderrey-Solana:2019ubu ; Wang:2019xey ; Li:2022tcr ; Budhraja:2023rgo . Measurements focused on the angular structure of inclusive jets indicate that jets become narrower in PbPb collisions at both the RHIC STAR:2021kjt and the LHC  ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , contrary to the theoretical expectation of intra-jet broadening Ringer:2019rfk ; Chien:2016led ; Milhano:2017nzm . In experiments, the medium modifications of the jet substructure are typically assessed by comparing two jet samples in PbPb and pp collisions that are selected within the same $p_{T}$ bins. Due to energy loss in the QGP, effectively quenched jets may be less likely to pass the $p_{T}$ selection threshold in A+A collisions, while some jets with insufficient quenching may still survive. This phenomenon is referred to as “selection bias” Baier:2001yt ; Renk:2012ve ; Wang:2021jgm ; Cunqueiro:2021wls ; Brewer:2021hmh ; Zhang:2021sua . Such biases could complicate the jet-by-jet comparison and obscure the connection between experimental measurements and intrinsic jet modifications Wang:2021jgm ; Brewer:2018dfs ; Du:2020pmp ; Brewer:2021hmh .
@@
The V+jet, jet tagged by the vector boson ($Z^{0}/W^{\pm}$ or $\gamma$), serves as a golden channel to explore the jet quenching phenomenon in high-energy heavy-ion collisions Neufeld:2010fj ; Wang:1996yh ; Dai:2012am ; Chen:2017zte ; Wang:2013cia ; Zhang:2018urd ; Chang:2019sae ; JETSCAPE:2024rma . Since the vector boson does not interact strongly with the hot nuclear matter, it effectively gauges the initial momentum of the recoiling jet. Furthermore, the V+jet process is dominated by quark-jet production, reducing the potential influences from changes in the $q/g$ fraction during A+A collisions Wang:2020qwe . By additionally constraining the $p_{T}$ of the vector boson, the impact of selection biases on jet measurements in A+A collisions was expected to be minimized Wang:2021jgm ; Brewer:2021hmh ; STAR:2023pal ; STAR:2023ksv . Therefore, V+jet may provide unique advantages to studying medium-induced jet broadening. Recently, the CMS collaboration reports the first measurement on the two angular structure observables of $\gamma$+jet, the jet girth ($g$Larkoski:2014pca and the groomed jet radius ($R_{g}$arkoski:2014wba , in pp and PbPb collisions at $\sqrt{s}=$ 5.02 TeV CMS:2024zjn . The results show that the medium modification pattern of the angular structure significantly depends on the selection cut, $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$, where $p_{T}^{\rm jet}$ and $p_{T}^{\rm\gamma}$ denote the transverse momentum of the jet and photon respectively. Notably, when setting $x_{j\gamma}>$ 0.4, there are hints of a broadening in jet angular structure at larger girth in PbPb collisions. This contrasts with the previously measured narrower girth distribution of inclusive jets reported by the ALICE collaboration ALICE:2018dxf ; ALargeIonColliderExperiment:2021mqf . The influences of selection bias in these two types of jet measurements warrant careful consideration. Timely theoretical explanations and quantitative investigations are essential for addressing this issue.
@@
This paper presents a theoretical study of the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. By employing a transport approach, we carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, which shows good agreement with the recently reported CMS data. Using the Jet-by-Jet matching method, we explore the relationship between selection bias and kinematic requirements in realistic event selection. Through quantitative analysis, we demonstrate that $\gamma$+jets can significantly reduce selection bias and effectively select sufficiently quenched jets in PbPb collisions compared to inclusive jets. We also discuss the influence of medium-induced gluon radiation and the medium response on the broadening of the jet angular substructure. For completeness, we will study the nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ distributions of both inclusive jets and $\gamma$+jets in PbPb collisions compared to the pp baseline.
@@
The remainder of this paper is organized as follows. In Sec. II, we will introduce the theoretical frameworks to study the angular structure of $\gamma$+jets in pp and PbPb collisions. In Sec. III, we will discuss the main results of this paper. At last, we will summarize this study in Sec. IV.
@@ II Theoretical framework
To investigate the angular structure of inclusive jets and $\gamma$+jets, we employ the PYTHIA8 Sjostrand:2014zea with the Monash Tune Skands:2014pea to generate the pp events as a baseline for the calculations of nucleus-nucleus collisions. Furthermore, we utilize a transport approach, which considers both the radiative and collisional partonic energy loss, to simulate the massive and massless jet evolution in the QGP. This hybrid transport approach has been used in the studies of light- and heavy-flavor dijet Dai:2018mhw ; Li:2024uzk ; Wang:2024yag ; Li:2024pfi , $Z^{0}/\gamma-$jet production in heavy-ion collisions Wang:2020qwe ; Wang:2023udp . Since the medium-induced gluon radiation plays a critical role in the jet energy loss Gyulassy:1993hr ; Wang:2001ifa , we use the radiation spectrum within the higher-twist formalism  Guo:2000nz ; Zhang:2003yn ; Zhang:2003wk ; Majumder:2009ge to simulate the in-medium jet shower in the hot/dense QCD matter,
@@
where $x$ and $k_{\perp}$ denote the energy fraction and transverse momentum carried by the radiated gluon. $\alpha_{s}$ is the strong coupling constant, $C_{s}$ the quadratic Casimir in color representation, $P(x)$ is the QCD splitting kernel Wang:2009qb for $q\rightarrow q+g$ and $g\rightarrow g+g$ respectively($g\rightarrow q+\bar{q}$ process is negligible due to its low possibility He:2011zx ),
@@
Moreover, $\tau_{f}=2Ex(1-x)/(k^{2}_{\perp}+x^{2}M^{2})$ is the gluon formation time considering the Landau-Pomeranchuk-Migdal (LPM) effects Wang:1994fx ; Zakharov:1996fv . $\hat{q}$ denotes the jet transport parameter Chen:2010te ; Rapp:2018qla ; Cao:2018ews , determined with a $\chi^{2}$ fitting to the identified hadron production in PbPb collisions at LHC Ma:2018swx . To consider the fluctuation of medium-induced gluon radiation, we assume that the number of the radiated gluon during a time step ($\Delta t=$0.1 fm) obeys the Poisson distribution $f(n)=\lambda^{n}e^{-\lambda}/{n!}$, where the parameter $\lambda$ denotes the mean number of the radiation calculated by integrating Eq. (1). Once the radiation number $n$ is determined, the corresponding energy-momentum can be further sampled by Eq. (1) one by one. More details of the Monte Carlo implementation of the medium-induced gluon radiation can be found in Refs. Cao:2016gvr ; Wang:2020qwe .
@@
In addition, it is also essential to consider the partonic energy loss from the elastic scattering. While the inelastic jet energy loss is carried out by the higher-twist formalisms, for completeness, the elastic energy loss is estimated by the pQCD calculation at the Hard Thermal Loop approximation Neufeld:2010xi ; Huang:2013vaa ,
$\frac{dE}{dL}=-\frac{\alpha_{s}C_{s}\mu_{D}^{2}}{2}{\rm ln}{\frac{\sqrt{ET}}{\mu_{D}}}.$ (4)
@@
where $L$ represents the transport path along to the parton’s momentum and $\mu_{D}^{2}=6\pi\alpha_{s}T^{2}$ the Debye screening mass. The collisional energy loss of a parton can be calculated by integrating Eq. (4) during each time step. This treatment is an adequate approximation since the medium-induced gluon radiation is the dominant energy loss mechanism for light parton. The initial spacial production vertex of jets in nucleus-nucleus collisions is determined based on the MC-Glauber model Miller:2007ri . In the simulation of a jet traversing the expanding fireball, we utilize the CLVisc hydrodynamic model Pang:2016igs to generate the temperature and velocity of each medium cell. When the local temperature reaches $T_{c}=0.165$ GeV, jet parton fragment into hadron with the Colorless Hadronization prescription, which the JETSCAPE collaboration developed Putschke:2019yrg based on the Lund string model Andersson:1983jt ; Sjostrand:1984ic .
@@
When studying jet substructures in high-energy nuclear collisions, it is important to consider the medium response effect. Energy transferred from high-$p_{T}$ jets can excite quasi-particles within the QGP medium KunnawalkamElayavalli:2017hxo ; Pablos:2019ngg ; Chen:2020tbl ; Casalderrey-Solana:2020rsj ; He:2018xjv ; Ke:2020clc . In this work, we utilize an approach based on the Cooper-Frye formula with perturbations Cooper:1974mv ; Casalderrey-Solana:2016jvj to take the medium response effect into account.
@@
×$\times\{p_{T}\Delta P_{T}{\rm cos(\Delta\phi)}+\frac{1}{3}m_{T}\Delta M_{T}{\rm cosh(\Delta y)}\}.$
Here, $\Delta y$ and $\Delta\phi$ represent the rapidity and azimuthal angle of the emitted thermal particles relative to the jet axis, respectively. Meanwhile, $m_{T}$ and $p_{T}$ denote their transverse mass and transverse momentum. The quantities $\Delta P_{T}$ and $\Delta M_{T}=\Delta E/\cosh y_{j}$ signify the transverse momentum and transverse mass transferred from the jet to the medium, where $\Delta E$ is the energy lost by the jets, and $T$ indicates the hadronization temperature of the emitted particles. Once the values for $\Delta P_{T}$ and $\Delta E$ of the jet during in-medium propagation are determined, one can sample the transverse momentum, rapidity, and azimuthal angle of the emitted particles individually based on Eq. (LABEL:eq:resp). For simplicity, we assume that all emitted partons from the excited medium hadronize into proton and pion ($\pi^{0},\pi^{\pm}$) with the ratio of $5\%$ for protons and $95\%$ for pions. Note that the above Eq. (LABEL:eq:resp) has not considered the radial flow of the expanding fireball, which cannot be neglected in the non-central nucleus-nucleus collisions. For this reason, we use the species- and momentum-dependent empirical expressions for the hadronization temperature $T$, as obtained by fitting the hadron spectra measured by the ALICE Casalderrey-Solana:2016jvj ; ALICE:2013mez .
$T_{\pi}({p_{\textrm{\tiny T}}})=\begin{cases}0.19~{}{\rm GeV},&{p_{\textrm{\tiny T}}}<0.7~{}{\rm GeV},\\ 0.21\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.28}\,\rm{GeV},&{p_{\textrm{\tiny T}}}>0.7~{}{\rm GeV},\end{cases}$ (6)
$T_{p}({p_{\textrm{\tiny T}}})=\begin{cases}0.15~{}{\rm GeV},&{p_{\textrm{\tiny T}}}<0.07~{}{\rm GeV},\\ 0.33\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.3}\rm{GeV},&0.07\sim 1.9~{}{\rm GeV},\\ 0.4~{}{\rm GeV},&{p_{\textrm{\tiny T}}}>1.9~{}{\rm GeV}.\end{cases}$ (7)

### v4

@@ Abstract
Medium modification of jet substructure in the hot and dense nuclear matter has garnered significant interest from the heavy-ion physics community in recent years. Measurements of inclusive jets show an angular narrowing in nucleus-nucleus collisions, while recent CMS results for photon-tagged jets ($\gamma$+jets) suggest evidence of broadening. In this study, we conduct a theoretical analysis of the angular structure of inclusive jets and $\gamma$+jets using a transport approach that accounts for jet energy loss and medium response in quark-gluon plasma. We examine the girth modification of $\gamma$+jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, achieving good agreement with recent CMS measurements. We explore the relationship between selection bias and jet kinematics by varying the threshold for $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. Notably, we quantitatively demonstrate that $\gamma$+jets significantly reduce selection bias and can effectively select jets that have been sufficiently quenched in PbPb collisions, which is crucial for capturing the jet angular broadening. Additionally, we estimate the contributions of medium-induced gluon radiation and medium response to the broadening of the jet angular substructure. Finally, we analyze the modification patterns of jet $R_{g}$ and $\Delta R_{\rm axis}$ in PbPb collisions, which indicate slight broadening for $\gamma$+jets and noticeable narrowing for inclusive jets compared to pp collisions.
@@ I Introduction
High-energy collisions of heavy nuclei at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) provide an experimental avenue to unravel the mysteries of quark-gluon plasma (QGP), a short-lived state of de-confined nuclear matter created at extremely high temperature and density. The jet quenching phenomenon, energy dissipation of an energetic parton when passing through hot and dense nuclear matter, is one of the most important signatures of the QGP formation Gyulassy et al. (2004); Gyulassy and Plumer (1990); Qin and Wang (2015); Vitev et al. (2008); Armesto et al. (2008); Vitev and Zhang (2010); Casalderrey-Solana et al. (2014); Gyulassy and Wang (1994); Wang and Guo (2001); Vitev and Zhang (2008). Investigations on jet quenching reveal the phase structure of the strongly-coupled nuclear matter and push our knowledge of the quantum chromodynamics (QCD) under extreme conditions Connors et al. (2018); Andrews et al. (2020); Cao and Wang (2021); Cunqueiro and Sickles (2022); Apolinário et al. (2022); Sorensen et al. (2024); Zhao et al. (2020); Xie et al. (2024); Zhang et al. (2022a); Tang et al. (2020); Shen and Yan (2020); Chen et al. (2025, 2024); Zhao et al. (2024); Shou et al. (2024).
@@
Jet substructures are valuable tools to gain insight into the details of the jet-medium interaction in QGP, such as the medium-induced gluon radiation Baier et al. (1997a, b), medium response Kunnawalkam Elayavalli and Zapp (2017); Pablos (2020); Chen et al. (2020); Casalderrey-Solana et al. (2021); He et al. (2019); Ke and Wang (2021), medium resolution length Hulcher et al. (2018); Mehtar-Tani and Tywoniuk (2017); Caucal et al. (2018) and the “Molière elastic scattering” D’Eramo et al. (2013, 2019). Recent reviews on this topic can be found in references Cunqueiro and Sickles (2022); Hayrapetyan et al. (2025a); Arslandok et al. (2023); Apolinário et al. (2022); Marzani et al. (2019). A significant focus of recent investigations has been on how the angular structure of jets is modified in nucleus-nucleus collisions—whether they narrow or broaden—an issue that has garnered much attention Ringer et al. (2020); Rajagopal et al. (2016); Chien and Vitev (2017); Mehtar-Tani and Tywoniuk (2017); Larkoski et al. (2017); Chang et al. (2018); Casalderrey-Solana et al. (2017); Kang et al. (2025a, b); Tachibana et al. (2024); Milhano et al. (2018); Caucal et al. (2019); Casalderrey-Solana et al. (2020); Wang et al. (2019); Li et al. (2023); Budhraja et al. (2025). Measurements focused on the angular structure of inclusive jets indicate that jets become narrower in PbPb collisions at both the RHIC Abdallah et al. (2022) and the LHC Acharya et al. (2022); Aad et al. (2023a); Acharya et al. (2018, 2023); Ehlers (2022); Aad et al. (2023b), contrary to the theoretical expectation of intra-jet broadening Ringer et al. (2020); Chien and Vitev (2017); Milhano et al. (2018). In experiments, the medium modifications of the jet substructure are typically assessed by comparing two jet samples in PbPb and pp collisions that are selected within the same $p_{T}$ bins. The sufficiently quenched jets, which traverse a longer path length and undergo relatively larger energy loss in the QGP, may be less likely to pass the $p_{T}$ selection threshold in AA collisions, while some jets with insufficient quenching may still survive. This phenomenon is referred to as “selection bias” Baier et al. (2001); Renk (2013); Wang et al. (2021); Cunqueiro and Sickles (2022); Brewer et al. (2022); Zhang et al. (2022b). Such biases could complicate the jet-by-jet comparison and obscure the connection between experimental measurements and intrinsic jet modifications Wang et al. (2021); Brewer et al. (2019); Du et al. (2020); Brewer et al. (2022).
@@
The V+jet, jet tagged by the vector boson ($Z^{0}/W^{\pm}$ or $\gamma$), serves as a golden channel to explore the jet quenching phenomenon in high-energy heavy-ion collisions Neufeld et al. (2011); Wang et al. (1996); Dai et al. (2013); Chen et al. (2018); Wang and Zhu (2013); Zhang et al. (2018); Chang et al. (2020); Sirimanna et al. (2024). Since the vector boson does not interact strongly with the hot nuclear matter, it effectively gauges the initial momentum of the recoiling jet. Furthermore, the V+jet process is dominated by quark-jet production, reducing the potential influences from changes in the $q/g$ fraction during AA collisions Wang et al. (2023). Additionally constraining the $p_{T}$ of the vector boson was expected to minimize the impact of selection biases on jet measurements in AA collisions Wang et al. (2021); Brewer et al. (2022); Aboona et al. (2025a, b). Therefore, V+jet may provide unique advantages to studying medium-induced jet broadening. Recently, the CMS collaboration reported the first measurement on the two angular structure observables of $\gamma$+jet, the jet girth (ggLarkoski et al. (2014a) and the groomed jet radius ($R_{g}$Larkoski et al. (2014b), in pp and PbPb collisions at $\sqrt{s}=5.02$ TeV Hayrapetyan et al. (2025b). The results show that the medium modification pattern of the angular structure significantly depends on the selection cut, $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$, where $p_{T}^{\rm jet}$ and $p_{T}^{\rm\gamma}$ denote the transverse momentum of the jet and photon respectively. Notably, when setting $x_{j\gamma}>$ 0.4, there are hints of a broadening in jet angular structure at larger girth in PbPb collisions. This contrasts with the previously measured narrower girth distribution of inclusive jets reported by the ALICE collaboration Acharya et al. (2018, 2022). The influences of selection bias in these two types of jet measurements warrant careful consideration. Timely theoretical explanations and quantitative investigations are essential for addressing this issue.
@@
This paper presents a theoretical study of the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. By employing a transport approach, we carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, which shows good agreement with recently reported CMS data. Using the Jet-by-Jet matching method, we explore the relationship between selection bias and kinematic requirements in realistic event selection. Through quantitative analysis, we demonstrate that $\gamma$+jets can significantly reduce selection bias and effectively select sufficiently quenched jets in PbPb collisions compared to inclusive jets. We also discuss the influence of medium-induced gluon radiation and medium response on the broadening of the jet angular substructure. For completeness, we will study the nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ distributions of both inclusive jets and $\gamma$+jets in PbPb collisions compared to the pp baseline.
@@
The remainder of this paper is organized as follows. In Section II, we will introduce the theoretical frameworks to study the angular structure of $\gamma$+jets in pp and PbPb collisions. In Section III, we will discuss the main results of this paper. At last, we will summarize this study in Section IV.
@@ II Theoretical framework
To investigate the angular structure of inclusive jets and $\gamma$+jets, we employ the PYTHIA8 Sjöstrand et al. (2015) with the Monash Tune Skands et al. (2014) to generate the pp events as a baseline for the calculations of nucleus-nucleus collisions. Furthermore, we utilize a transport approach, which considers both the radiative and collisional partonic energy loss, to simulate massive and massless jet evolution in QGP. This hybrid transport approach has been used in the studies of light- and heavy-flavor dijet Dai et al. (2020); Li et al. (2024); Wang et al. (2025a); Li et al. (2025), $Z^{0}/\gamma+$jet production in heavy-ion collisions Wang et al. (2023, 2025b). Since the medium-induced gluon radiation plays a critical role in jet energy loss Gyulassy and Wang (1994); Wang and Guo (2001), we use the radiation spectrum within the higher-twist formalism Guo and Wang (2000); Zhang and Wang (2003); Zhang et al. (2004); Majumder (2012) to simulate the in-medium jet shower in the hot/dense QCD matter,
@@
where $x$ and $k_{\perp}$ denote the energy fraction and transverse momentum carried by the radiated gluon. $\alpha_{s}$ is the strong coupling constant, $C_{s}$ the quadratic Casimir in color representation, $P(x)$ is the QCD splitting kernel Deng and Wang (2010) for $q\rightarrow q+g$ and $g\rightarrow g+g$ respectively($g\rightarrow q+\bar{q}$ process is negligible due to its low possibility He et al. (2012)),
@@
Moreover, $\tau_{f}=2Ex(1-x)/(k^{2}_{\perp}+x^{2}M^{2})$ is the gluon formation time considering the Landau-Pomeranchuk-Migdal (LPM) effect Wang et al. (1995); Zakharov (1996). $\hat{q}=q_{0}(T/T_{0})^{3}$ denotes the jet transport parameter Chen et al. (2010); Beraudo et al. (2018); Cao et al. (2019), where $T_{0}$ is the medium temperature at the center of QGP at $\tau=0.6$ fm. In this work, we use the extracted value of $q_{0}=1.2$ GeV2/fm determined with a $\chi^{2}$ fitting to the identified hadron production in PbPb collisions at LHC Ma et al. (2019). To consider the fluctuation of medium-induced gluon radiation, we assume that the number of the radiated gluon during a time step ($\Delta t=0.1$ fm) obeys the Poisson distribution $f(n)=\lambda^{n}e^{-\lambda}/{n!}$, where the parameter $\lambda$ denotes the mean number of the radiation calculated by integrating Eq. (1). Once the radiation number $n$ is determined, the corresponding energy-momentum can be further sampled by Eq. (1) one by one. Due to the LPM effect, a radiated gluon can only interact independently with the medium after a formation time $\tau_{f}$ and then further lose energy (including medium-induced gluon radiation). Note that Eq. (1) represents the net radiative contribution at the high-energy limit as derived from the complete twist-4 calculations Chen et al. (2010); Guo and Wang (2000); Zhang and Wang (2003); Zhang et al. (2004). We impose a lower energy cutoff for the radiated gluons at the Debye screening mass $\mu_{D}$ to avoid the infrared behavior of Eq. (1) in the limit $x\rightarrow 0$, where $\mu_{D}^{2}=4\pi\alpha_{s}(1+N_{f}/6)T^{2}$ with $N_{f}=3$.
@@
In addition, it is also essential to consider the partonic energy loss from the elastic scattering. While the inelastic jet energy loss is described using higher-twist formalisms, for completeness, the elastic energy loss is estimated via pQCD calculations within the Hard-Thermal-Loop (HTL) approximation at leading logarithmic accuracy Braaten and Thoma (1991); Neufeld (2011),
$\frac{dE}{dL}=-\frac{C_{s}\alpha_{s}g_{s}^{2}T^{2}}{2}(1+\frac{N_{f}}{6}){\rm ln}{\frac{q_{\rm max}}{q_{\rm min}}},$
@@
where $L$ represents the transport path along the parton’s momentum direction. We set physically reasonable upper and lower limits for the 3-momentum transfer $q$ as $q_{\rm max}\sim\sqrt{ET}$ and $q_{\rm min}\sim m_{D}$. The collisional energy loss of a parton can be calculated by integrating Eq. (II) during each time step $\Delta t$ according to the parton energy and the temperature of the medium cell. This treatment is an adequate approximation since the medium-induced gluon radiation is the dominant energy loss mechanism for light partons. The initial spatial production vertex of jets in nucleus-nucleus collisions is determined based on the MC-Glauber model Miller et al. (2007). In the simulation of a jet traversing the expanding fireball, we utilize the CLVisc hydrodynamic model Pang et al. (2016) to generate the temperature and velocity of each medium cell. When the local temperature reaches $T_{c}=0.165$ GeV, jet partons fragment into hadrons with the Colorless Hadronization prescription, which the JETSCAPE collaboration developed Putschke et al. (2019) based on the Lund string model Andersson et al. (1983); Sjostrand (1984).
@@
When studying jet substructures in high-energy nuclear collisions, it is important to consider the medium response effect. Energy transferred from high-$p_{T}$ jets can excite quasi-particles within the QGP medium Kunnawalkam Elayavalli and Zapp (2017); Pablos (2020); Chen et al. (2020); Casalderrey-Solana et al. (2021); He et al. (2019); Ke and Wang (2021). In this work, we utilize an approach based on the Cooper-Frye formula with perturbations Cooper and Frye (1974); Casalderrey-Solana et al. (2017) to take the medium response effect into account.
@@
$\times\{p_{T}\Delta P_{T}{\rm cos(\Delta\phi)}+\frac{1}{3}m_{T}\Delta M_{T}{\rm cosh(\Delta y)}\}.$
Here, $\Delta y$ and $\Delta\phi$ represent the rapidity and azimuthal angle of the emitted thermal particles relative to the jet axis, respectively. Meanwhile, $m_{T}$ and $p_{T}$ denote their transverse mass and transverse momentum. The quantities $\Delta P_{T}$ and $\Delta M_{T}=\Delta E/\cosh y_{j}$ signify the transverse momentum and transverse mass transferred from the jet to the medium, where $\Delta E$ is the energy lost by the jets, and $T$ indicates the hadronization temperature of the emitted particles. Once the values for $\Delta P_{T}$ and $\Delta E$ of the jet during in-medium propagation are determined, one can sample the transverse momentum, rapidity, and azimuthal angle of the emitted particles individually based on Eq. (LABEL:eq:resp). All emitted partons from the excited medium are assumed to hadronize into protons and pions ($\pi^{0},\pi^{\pm}$) with a ratio of $5\%$ for protons and $95\%$ for pions. This branching ratio was approximately extracted in the Hybrid model Casalderrey-Solana et al. (2017, 2021) by fitting the particle spectra from PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV measured by the ALICE collaboration Abelev et al. (2013). In that work, the hadronization temperature $T$ was parameterized using the following species- and momentum-dependent empirical expressions.
$\hskip-21.68121ptT_{\pi}({p_{\textrm{\tiny T}}})=\begin{cases}0.19~{\rm GeV},&{p_{\textrm{\tiny T}}}<0.7~{\rm GeV}\\ 0.21\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.28}\,\rm{GeV},&{p_{\textrm{\tiny T}}}>0.7~{\rm GeV}\end{cases}$ (5)
$\hskip-21.68121ptT_{p}({p_{\textrm{\tiny T}}})=\begin{cases}0.15~{\rm GeV},&{p_{\textrm{\tiny T}}}<0.07~{\rm GeV}\\ 0.33\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.3}\rm{GeV},&0.07\sim 1.9~{\rm GeV}\\ 0.4~{\rm GeV},&{p_{\textrm{\tiny T}}}>1.9~{\rm GeV}\end{cases}$ (6)